: CPI baske (15 tbs beef, 5 ths chicken price of heef price of chicken s15 s10 Suppose 2004 is the base year a. Compute the CPl in 2005 b. What was the CPI inflation rate from 2005-20067 e. 2004-5: Households bought CPI basket beef, 15 ths chicken). In 2006: Households bought (5 Ibs Compute cost of the 2006 household bask et, And compute % increase in cost of household basket over 2005-6
Answer #1

a) CPI = Cost of Market basket in given year / Cost of Market basket in base year * 100

= 8*15 + 8*5 / 6*15 + 6*5 * 100

= 120 + 40 / 90 + 30 * 100

= 160 / 120 * 100

= 133.33

b) CPI in 2005 = 15*8 + 5*8 / 15*6 + 5*6 * 100

120 + 40 / 90 + 30 * 100

160 / 120 * 100

= 133.33

CPI in 2006 = 15*15 + 10*5 / 15*8 + 5*8 * 100

= 225 + 50 / 120 + 40 * 100

= 275 / 160 * 100

= 171.88

CPI Inflation = 171.88 - 133.33 / 133.33 * 100

= 28.91%

c) Cost of 2006 household basket

= 15*5 + 15*10

= 75 + 150

= 225

% Increase in cost of household basket 2005-06

Cost in 2005 = 15*8 + 5*8

= 120 + 40

= 160

% Increase = 225 - 160 / 160 * 100

= 108.33 %